# Complementarity-based complementarity

### v2

@@ III Results and discussion
For the experimental verification of the bounds (1) and (2), we probed different types of input states and, for each state, we calculated the entropy sum $H(\hat{X})+H(\hat{Y})+H(\hat{Z})$ from the measured output probabilities. The results for the two MUB triplets $CDF$ and $ABE$ are shown as filled markers in Figure 5, and compared to their respective lower entropy bounds (1) (green dashed line) and (2) (red dash-dotted line). For each input state, the error bars are calculated by sampling 500 sets of counts from a distribution with the measured mean and standard deviation of the original dataset and taking the 10%-90% spread of the corresponding entropy distribution. The error bars are not visible in most points due to their narrow extent.
@@
The figure also shows the entropy values predicted via realistic simulations of the measurement process. These simulations are obtained by first characterizing the performance of the mQPG-based decoder through a quantum detector tomography [22, 23, 19] to reconstruct the actual POVMs. The average measurement error (cross-talk) per basis falls between 0.5% (for basis $|a_{i}\rangle$) and 1.9% (for basis $|f_{n}\rangle$). From the reconstructed POVMs, we calculated the expected entropy that one would observe measuring each state with this imperfect system, obtaining the hollow markers in Fig. 5. These estimates almost perfectly match the measured values, confirming that discrepancies with the theoretical predictions are to be attributed to the imperfect detection system.
@@
The first type of probed input states are the eigenstates of all six MUBs, labelled “internal” if they are eigenstates of $\hat{X}$, $\hat{Y}$ or $\hat{Z}$, and “external” otherwise. The entropy sum of the internal eigenstates (green squares) is always equal to the value predicted in (1) which, for the triplet $CDF$, is the minimum possible entropy sum. Contrarily, the external eigenstates (blue diamonds) always maximize the entropy sum.
@@
Then we probed random input states (orange circles), generated by sampling amplitude and phase coefficients from a uniform distribution and renormalizing the amplitudes. We note that this method does not allow for truly uniform sampling of the parameter space; however, this is not relevant in the scope of this work, as we only look at random input states to verify that they fall within the predicted entropy boundaries.
@@
Finally, we probed states that are a superposition of four states of the computational basis $|a_{i}\rangle$, e.g. $\sum_{n}\psi_{n}e^{i\phi_{n}}|a_{n}\rangle$ with $\{\psi_{n}\}=(0.19323,0.68019,0,0.68019,0.19323)$ and $\{\phi_{n}\}=\left(-{3}\pi/{5},{\pi}/{5},0,0,0\right)$ (red triangles in Fig. 5). These states were initially proven to minimize the sum of entropies in four, five and six MUBs in $d=5$ [9]. However, while testing this assumption, we observed that the sum of entropies in the $ABE$ triplet violated bound (1), previously assumed to hold for any possible choice of MUB triplet. This contradiction led to the discovery of bound (2) for the basis triplet $ABE$, while $CDF$ maintains the known bound, revealing the underlying asymmetry between sets of MUBs.
@@ IV Conclusions
In conclusion we have shown that the amount of joint information that one can have on maximally complementary observables can depend on which of them are considered, even though the overlap between the eigenstates of all maximally complementary observables are all the same, since they are MUBs. We showed that this holds in dimension $d=5$ when the maximum joint information is gauged by minimizing the sum of the entropies and of the variances of three MUBs. For smaller dimensions, these metrics do not show this effect.
@@ Appendix A MUBs, explicit form
$\{|a_{i}\rangle\}=$ $\begin{pmatrix}1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\\ 0&0&0&1&0\\ 0&0&0&0&1\end{pmatrix}$ (6)
$\{|b_{j}\rangle\}=\frac{1}{\sqrt{5}}$ $\begin{pmatrix}1&1&1&1&1\\ 1&\omega&\omega^{2}&\omega^{3}&\omega^{4}\\ 1&\omega^{2}&\omega^{4}&\omega&\omega^{3}\\ 1&\omega^{3}&\omega&\omega^{4}&\omega^{2}\\ 1&\omega^{4}&\omega^{3}&\omega^{2}&\omega\end{pmatrix}$
$\{|c_{k}\rangle\}=\frac{1}{\sqrt{5}}$ $\begin{pmatrix}1&1&1&1&1\\ \omega&\omega^{2}&\omega^{3}&\omega^{4}&1\\ \omega^{4}&\omega&\omega^{3}&1&\omega^{2}\\ \omega^{4}&\omega^{2}&1&\omega^{3}&\omega\\ \omega&1&\omega^{4}&\omega^{3}&\omega^{2}\end{pmatrix}$
$\{|d_{l}\rangle\}=\frac{1}{\sqrt{5}}$ $\begin{pmatrix}1&1&1&1&1\\ \omega^{3}&\omega^{4}&1&\omega&\omega^{2}\\ \omega^{2}&\omega^{4}&\omega&\omega^{3}&1\\ \omega^{2}&1&\omega^{3}&\omega&\omega^{4}\\ \omega^{3}&\omega^{2}&\omega&1&\omega^{4}\end{pmatrix}$
$\{|e_{m}\rangle\}=\frac{1}{\sqrt{5}}$ $\begin{pmatrix}1&1&1&1&1\\ \omega^{2}&\omega^{3}&\omega^{4}&1&\omega\\ \omega^{3}&1&\omega^{2}&\omega^{4}&\omega\\ \omega^{3}&\omega&\omega^{4}&\omega^{2}&1\\ \omega^{2}&\omega&1&\omega^{4}&\omega^{3}\end{pmatrix}$
$\{|f_{n}\rangle\}=\frac{1}{\sqrt{5}}$ $\begin{pmatrix}1&1&1&1&1\\ \omega^{4}&1&\omega&\omega^{2}&\omega^{3}\\ \omega&\omega^{3}&1&\omega^{2}&\omega^{4}\\ \omega&\omega^{4}&\omega^{2}&1&\omega^{3}\\ \omega^{4}&\omega^{3}&\omega^{2}&\omega&1\end{pmatrix},$
@@
$\{|a_{i}\rangle\}=$ $\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\end{pmatrix}\,\,\,\{|b_{j}\rangle\}=\frac{1}{2}\begin{pmatrix}1&1&1&1\\ 1&1&-1&-1\\ 1&-1&-1&1\\ 1&-1&1&-1\end{pmatrix}$ (7)
$\{|c_{k}\rangle\}=\frac{1}{2}$ $\begin{pmatrix}1&1&1&1\\ 1&1&-1&-1\\ -i&i&i&-i\\ i&-i&i&-i\end{pmatrix}$
$\{|d_{l}\rangle\}=\frac{1}{2}$ $\begin{pmatrix}1&1&1&1\\ i&-i&i&-i\\ -1&-1&1&1\\ i&-i&-i&i\\ \end{pmatrix}$
$\{|e_{m}\rangle\}=\frac{1}{2}$ $\begin{pmatrix}1&1&1&1\\ i&-i&i&-i\\ i&-i&-i&i\\ -1&-1&1&1\end{pmatrix}.$
@@ Appendix B Numerical minimization
i.e. a parametrization yielding the normalization of the state intrinsically. Differently from Ref. [9], we checked the results of our optimizations by repeating the procedure twice: The first one, on re-parametrized states where the weights in Eq. (8) were permuted with respect to the states of the basis, and the second one on each of the re-paramterized states but expanded on the eigenstates of other MUBs.
For each MUB $\{|x_{k}\rangle\}$ in Eq. (6) we retrieved the Born probabilities in terms of the coefficients in Eq. (8) through the superpositions between the states of the MUB and the generic state $|\psi\rangle$, namely
@@
with $\hat{X}=\sum_{k}P(k)|x_{k}\rangle\langle x_{k}|$. We then minimized numerically the sum of three entropies and the sum of three variances for all possible triplets of distinct MUBs. In the case of the variances, we also minimized over every permutation $P(k)$ of the eigenvalues $k$. The minimization was performed by using the routine NMinimize of Wolfram Mathematica, which finds the minimum of a function over a given set of parameters and constraints.

### v4

@@ III Results and discussion
For the experimental verification of the bounds in Eqs. (1) and (2), we probed different types of input states and, for each state, we calculated the entropy sum H(X^)+H(Y^)+H(Z^)H($\hat${X})+H($\hat${Y})+H($\hat${Z}) from the measured output probabilities. The results for the two MUB triplets $CDF$ and $ABE$ are shown as filled markers in Figure 5, and compared to their respective lower entropy bounds from Eqs. (1) (green dashed line) and (2) (red dash-dotted line). For each input state, the error bars are calculated by sampling 500 sets of counts from a distribution with the measured mean and standard deviation of the original dataset and taking the 10%-90% spread of the corresponding entropy distribution. The error bars are not visible in most points due to their narrow extent.
@@
The figure also shows the entropy values predicted via realistic simulations of the measurement process. These simulations are obtained by first characterizing the performance of the mQPG-based decoder through a quantum detector tomography [28, 29, 25] to reconstruct the actual POVMs. The average measurement error (cross-talk) per basis falls between 0.5% (for basis $\{|a_{j}\rangle\}_{j}$) and 1.9% (for basis $\{|f_{j}\rangle\}_{j}$). From the reconstructed POVMs, we calculated the expected entropy that one would observe measuring each state with this imperfect system, obtaining the hollow markers in Fig. 5. These estimates almost perfectly match the measured values, confirming that discrepancies with the theoretical predictions are to be attributed to the imperfect detection system.
@@
The first type of probed input states are the eigenstates of all six MUBs, labelled “internal” if they are eigenstates of $\hat{X}$, $\hat{Y}$ or $\hat{Z}$, and “external” otherwise. The entropy sum of the internal eigenstates (green squares) is always equal to the value predicted in Eq. (1) which, for the triplet $CDF$, is the minimum possible entropy sum. Contrarily, the external eigenstates (blue diamonds) always maximize the entropy sum.
@@
Then we probed pure random input states (orange circles), generated by sampling amplitude and phase coefficients from a uniform distribution and renormalizing the amplitudes. We note that this method does not allow for truly uniform sampling of the parameter space; however, this is not relevant in the scope of this work, as we only look at random input states to verify that they fall within the predicted entropy boundaries.
@@
Finally, we probed states that are a superposition of four states of the computational basis $|a_{i}\rangle$, e.g. $\sum_{j}\psi_{j}e^{i\phi_{j}}|a_{j}\rangle$ with $\{\psi_{j}\}=(0.19323,0.68019,0,0.68019,0.19323)$ and $\{\phi_{j}\}=\left(-{3}\pi/{5},{\pi}/{5},0,0,0\right)$ (red triangles in Fig. 5). These states were initially proven to minimize the sum of entropies in four, five and six MUBs in $d=5$ [17]. However, while testing this assumption, we observed that the sum of entropies in the $ABE$ triplet violated bound (1), previously assumed to hold for any possible choice of MUB triplet. This contradiction led to the discovery of bound in Eq. (2) for the basis triplet $ABE$, while $CDF$ maintains the known bound, revealing the immediate experimental consequences of the underlying asymmetry between sets of inequivalent MUBs.
Interestingly, while our primary goal in this demonstration was to highlight the effect of inequivalent sets of MUBs on uncertainty relations, our findings also provide a new experimental verification of their existence that can, in principle, be extended to higher dimensions. In fact, while the existence and implications of inequivalent classes of MUBs have been theoretically understood, their experimental verification has remained elusive until very recently due to the subtlety of these effects, which could only be observed by a setup capable of performing simultaneous projections across a complete and arbitrary basis. Ref. [30] provided an important verification in $d=4$ using hybrid path-polarization encodings; however, their method is based on indirect estimations, and lacks straightforward scalability to higher dimensions. In contrast, our demonstration relates the existence of inequivalent sets of MUBs to the simplest and most direct type of measurements. The versatility of the mQPG operating in the time-frequency degree of freedom makes our approach naturally extensible [25, 31], enabling experimental verification of similar UR bounds if they are discovered in higher-dimensional systems.
@@ IV Conclusions
In conclusion we have shown that the URs are shaped by the inequivalence of sets of MUBs, which is a physical fundamental proof that the amount of joint information that one can have on maximally complementary observables can depend on which of them are considered, even though the overlap between the eigenstates of all maximally complementary observables are all the same, since they are MUBs. We showed that two distinct URs appear for triplets of MUBs in dimension $d=5$, when the maximum joint information is gauged by minimizing the sum of the entropies and of the variances.
This effect has several implications which will be explored in future work. For instance, in quantum key distribution, the entropic uncertainty relations are known to provide a tight bound on the secret key rate and are crucial for security proofs. In addition, high-dimensional systems provide better key rates and larger maximum tolerable errors. Then, complementarity-based complementarity may be exploited to add a further level of security. Moreover, uncertainty relations are at the basis of many quantum information procedures (e.g. quantum metrology). Another implication refers to the foundations of quantum mechanics, where the investigation of URs in large dimensional systems can shed light on their complementary properties.
@@ Appendix A MUBs, explicit form
$\{|a_{j}\rangle\}_{j=1}^{5}=$ $\begin{pmatrix}1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\\ 0&0&0&1&0\\ 0&0&0&0&1\end{pmatrix}$ (8)
$\{|b_{j}\rangle\}_{j=1}^{5}=\frac{1}{\sqrt{5}}$ $\begin{pmatrix}1&1&1&1&1\\ 1&\omega&\omega^{2}&\omega^{3}&\omega^{4}\\ 1&\omega^{2}&\omega^{4}&\omega&\omega^{3}\\ 1&\omega^{3}&\omega&\omega^{4}&\omega^{2}\\ 1&\omega^{4}&\omega^{3}&\omega^{2}&\omega\end{pmatrix}$
$\{|c_{j}\rangle\}_{j=1}^{5}=\frac{1}{\sqrt{5}}$ $\begin{pmatrix}1&1&1&1&1\\ \omega&\omega^{2}&\omega^{3}&\omega^{4}&1\\ \omega^{4}&\omega&\omega^{3}&1&\omega^{2}\\ \omega^{4}&\omega^{2}&1&\omega^{3}&\omega\\ \omega&1&\omega^{4}&\omega^{3}&\omega^{2}\end{pmatrix}$
$\{|d_{j}\rangle\}_{j=1}^{5}=\frac{1}{\sqrt{5}}$ $\begin{pmatrix}1&1&1&1&1\\ \omega^{3}&\omega^{4}&1&\omega&\omega^{2}\\ \omega^{2}&\omega^{4}&\omega&\omega^{3}&1\\ \omega^{2}&1&\omega^{3}&\omega&\omega^{4}\\ \omega^{3}&\omega^{2}&\omega&1&\omega^{4}\end{pmatrix}$
$\{|e_{j}\rangle\}_{j=1}^{5}=\frac{1}{\sqrt{5}}$ $\begin{pmatrix}1&1&1&1&1\\ \omega^{2}&\omega^{3}&\omega^{4}&1&\omega\\ \omega^{3}&1&\omega^{2}&\omega^{4}&\omega\\ \omega^{3}&\omega&\omega^{4}&\omega^{2}&1\\ \omega^{2}&\omega&1&\omega^{4}&\omega^{3}\end{pmatrix}$
$\{|f_{j}\rangle\}_{j=1}^{5}=\frac{1}{\sqrt{5}}$ $\begin{pmatrix}1&1&1&1&1\\ \omega^{4}&1&\omega&\omega^{2}&\omega^{3}\\ \omega&\omega^{3}&1&\omega^{2}&\omega^{4}\\ \omega&\omega^{4}&\omega^{2}&1&\omega^{3}\\ \omega^{4}&\omega^{3}&\omega^{2}&\omega&1\end{pmatrix}$,
@@
$\{|a_{j}\rangle\}_{j=1}^{4}=$ $\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\end{pmatrix}$ (9)
$\{|b_{j}\rangle\}_{j=1}^{4}=\frac{1}{2}$ $\begin{pmatrix}1&1&1&1\\ 1&1&-1&-1\\ 1&-1&-1&1\\ 1&-1&1&-1\end{pmatrix}$
$\{|c_{j}\rangle\}_{j=1}^{4}=\frac{1}{2}$ $\begin{pmatrix}1&1&1&1\\ 1&1&-1&-1\\ -i&i&i&-i\\ i&-i&i&-i\end{pmatrix}$
$\{|d_{j}\rangle\}_{j=1}^{4}=\frac{1}{2}$ $\begin{pmatrix}1&1&1&1\\ i&-i&i&-i\\ -1&-1&1&1\\ i&-i&-i&i\\ \end{pmatrix}$
$\{|e_{j}\rangle\}_{j=1}^{4}=\frac{1}{2}$ $\begin{pmatrix}1&1&1&1\\ i&-i&i&-i\\ i&-i&-i&i\\ -1&-1&1&1\end{pmatrix}$.
@@ Appendix B Numerical minimization
i.e. a parametrization yielding the normalization of the state intrinsically. Differently from Ref. [17], we checked the results of our optimizations by repeating the procedure twice: The first one, on re-parametrized states where the weights in Eq. (10) were permuted with respect to the states of the basis, and the second one on each of the re-parametrized states but expanded on the eigenstates of other MUBs.
For each MUB $\{|x_{k}\rangle\}_{k}$ in Eq. (8) we retrieved the Born probabilities in terms of the coefficients in Eq. (10) through the superpositions between the states of the MUB and the generic state $|\psi\rangle$, namely
@@
with $\hat{X}=\sum_{k}P(k)|x_{k}\rangle\langle x_{k}|$. We then minimized numerically the sum of three entropies and the sum of three variances for all possible triplets of distinct MUBs. The minimization was performed by using the routine NMinimize of Wolfram Mathematica, which finds the minimum of a function over a given set of parameters and constraints.
Appendix C Effects of the inequivalence of MUBs on the EURs
Here we inspect more in detail the effect of the inequivalence of triplets of MUBs on the pertaining EURs. In particular, we show that the combination of the MUBs in the triplets from $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ identifies two related subsets of the Born probabilities inspected so far.
The MUBs shown in the appendix are Hadamard matrices related to each other through the unitary $\hat{U}={\rm diag}(1,\omega,\omega^{4},\omega^{4},\omega)$, with $\omega=\exp{(i2\pi/5)}$, as follows
[…]
$\{|e_{j}\rangle\}=\hat{U}^{2}\{|b_{j}\rangle\}$
$\{|d_{j}\rangle\}=\hat{U}^{3}\{|b_{j}\rangle\}$
$\{|f_{j}\rangle\}=\hat{U}^{4}\{|b_{j}\rangle\}.$
This structure inherently prevents from mapping any triplet in the set $\mathcal{S}_{1}$ of Eq. (I) into a triplet in $\mathcal{S}_{2}$ of Eq. (I) by means of a unitary matrix [8]. In the following Lemma, we show how these relations impact on the Born probabilities appearing in the URs. In particular, the proof outlines how the structure of the five-dimensional MUBs implies the same EUR in Eq. (1) (Eq. (2)) for the triplets in the same set $\mathcal{S}_{1}$ ($\mathcal{S}_{2}$).
Lemma 1.
Let $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ be the two inequivalent subsets of triplets of MUBs in $d=5$. Be
$\mathcal{P}_{1}\equiv\{p_{X_{1},\psi_{0}}^{(j)},p_{Y_{1},\psi_{0}}^{(j^{\prime})},p_{Z_{1},\psi_{0}}^{(j^{\prime\prime})}\}_{j,j^{\prime},j^{\prime\prime}}$ (14)
and
$\mathcal{P}_{2}\equiv\{p_{X_{2},\psi_{0}}^{(j)},p_{Y_{2},\psi_{0}}^{(j^{\prime})},p_{Z_{2},\psi_{0}}^{(j^{\prime\prime})}\}_{j,j^{\prime},j^{\prime\prime}}$ (15)
the two disjoint sets of probability vectors with $X_{1}Y_{1}Z_{1}\in\mathcal{S}_{1}$ and $X_{2}Y_{2}Z_{2}\in\mathcal{S}_{2}$, where $p_{T,\psi_{0}}^{(j)}$ is the Born probability of the superposition of the $j$-th eigenstate of a basis $T\in\{A,B,C,D,E,F\}$ with a five-dimensional pure state $|\psi_{0}\rangle$ and $\{p_{T,\psi_{0}}^{(j)}\}_{j=1}^{5}$ is the pertaining probability vector. Consider the unitary $U={\rm diag}(1,\omega,\omega^{4},\omega^{4},\omega)$, with $\omega=\exp{(i2\pi/5)}$, and the Fourier matrix $\Phi\equiv\{b_{j}\}_{j}$. Then the transformations in the set $\mathcal{V}\equiv\{U^{N}\Phi^{M}U^{L}\}_{N,L,M}$, with $N,L\in\{0,\pm 1\mod{5},\pm 2\mod{5}\}$ and $M\in\{0,\pm 1\}$, are authomorphisms of $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$.
Proof.
The probability $p_{T,\psi_{0}}^{(j)}$ reads
$p_{T,\psi_{0}}^{(j)}=|\langle\psi_{0}|t_{j}\rangle|^{2}$ (16)
where $\{|t_{j}\rangle\}_{j}$ are the eigenstates of $T$, identifying one of the MUBs. Due to the construction of the MUBs in terms of Hadamard matrices, the probabilities $p_{T,\psi_{0}}^{(j)}$ can be always be expressed as
$p_{T,\psi_{0}}^{(j)}=|\langle\psi_{0}|U^{n_{T}}\Phi^{m_{T}}U^{l}P_{jk}|a_{k}\rangle|^{2}$ (17)
where $\{|a_{j}\rangle\}_{j}$ is the computational basis and $P_{jk}$ is the permutation exchanging the vector $|a_{k}\rangle$ with $|a_{j}\rangle$. Being $U$ a periodic matrix such that $U^{5}=I$, we have $l,n_{T}\in\{0,\pm 1\mod{5},\pm 2\mod{5}\}$, while, for the Fourier matrix, we just need $m_{T}=0,\pm 1$. Permutations map the probability vectors $\{p_{T,\psi_{0}}^{(j)}\}_{j}$ in themselves and are therefore automorphisms for $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$. Similarly, the action of $\hat{U}^{l}$ on the states of the computational basis reduces to the multiplication for a global phase factor, thus leaving $p_{T,\psi_{0}}^{(j)}$ unchanged. Conversely, the exponents $m_{T}$ and $n_{T}$ determines the basis $T$. If $m_{T}=0$, the transformation $U^{n_{T}}\Phi^{m_{T}}$, again, rescales $|a_{j}\rangle$ by an irrelevant phase factor, while, for $m_{T}=\pm 1$, maps it to the a state of another MUB, identified by $n_{T}$. Now, we evaluate the probabilities in Eq. (17) on a state $|\psi\rangle=V|\psi_{0}\rangle$, with unitary $V$. By definition of $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$, if $V^{\dagger}$ is an automorphism of the set of MUBs, namely $V^{\dagger}:T\rightarrow\overline{T}$, with $T$ and $\overline{T}$ mutually unbiased, as
$V^{\dagger}U^{n_{T}}\Phi^{m_{T}}={U^{n_{\overline{T}}}}{\Phi^{m_{\overline{T}}}},$ (18)
then it is an automorphism of $\mathcal{P}_{1}\cup\mathcal{P}_{2}$, and
$V^{\dagger}={U^{n_{\overline{T}}}}{\Phi^{m_{\overline{T}}-m_{T}}}{U^{n_{T}}}\equiv U^{N}\Phi^{M}U^{L}\in\mathcal{V}.$ (19)
Now we have to show that the transformations in $\mathcal{V}$ identify the two disjoint sets $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$. The application of the same unitary $V$ to the state $|\psi_{0}\rangle$ relates the triplets of probability vectors $\{p_{X,\psi}^{(j)},p_{Y,\psi}^{(j^{\prime})},p_{Z,\psi}^{(j^{\prime\prime})}\}$ and $\{p_{\overline{X},\psi_{0}}^{(j)},p_{\overline{Y},\psi_{0}}^{(j^{\prime})},p_{\overline{Z},\psi_{0}}^{(j^{\prime\prime})}\}$ through the equations
$U^{N}\Phi^{M}U^{L+n_{X}}\Phi^{m_{X}}=U^{n_{\overline{X}}}\Phi^{m_{\overline{X}}}$ (20)
$U^{N}\Phi^{M}U^{L+n_{Y}}\Phi^{m_{Y}}=U^{n_{\overline{Y}}}\Phi^{m_{\overline{Y}}}$ (21)
$U^{N}\Phi^{M}U^{L+n_{Z}}\Phi^{m_{Z}}=U^{n_{\overline{Z}}}\Phi^{m_{\overline{Z}}}$ (22)
Without loss of generality, we can restrict to the cases where $M=0,L\neq 0$ and $M\neq 0,L=0$, being the remaining cases retrieved by shifting $n_{X}$, $n_{Y}$ and $n_{Z}$ by $L$. For the sake of simplicity, we will set the triplet $XYZ$ in the so called standard form, where one of the MUBs is the computational basis, say $X=\{|a_{j}\rangle\}_{j}$, implying $m_{X}=0$. The proof with $m_{X}=\pm 1$ follows the same line of argument and leads to the same conclusions.
In the case $M=0$, note that we can also set $N=0$. Then, Eqs. (20), (21) and (22) fix $m_{\overline{X}}=m_{X}=0$, $m_{\overline{Y}}=m_{Y}\neq 0$ and $m_{\overline{Z}}=m_{Z}\neq 0$, respectively. Note that the last two must be both non-null, or we would have repetitions of the computational basis in the same triplet. Finally, Eqs. (21) and (22) set $\Delta n\equiv n_{Z}-n_{Y}=n_{\overline{Z}}-n_{\overline{Y}}$, implying that the application of $V$ fixes a relation between $\{Y,Z\}$ and $\{\overline{Y},\overline{Z}\}$. From the periodicity of $U$, again $\Delta n\in\{0,\pm 1\mod{5},\pm 2\mod{5}\}$. We must have $\Delta n\neq 0$ to avoid $Y$ and $Z$ being the same basis. Moreover, the order in the difference $\Delta n$ is irrelevant, and so is its sign. These considerations leave us with two distinct possibilities, either $\Delta n=\pm 1\mod{5}$ or $\Delta n=\pm 2\mod{5}$. But these identify two disjoint sets of triplets, the one in $\mathcal{S}_{1}$, featuring $\Delta n=\pm 1\mod{5}$, and the one in $\mathcal{S}_{2}$, with $\Delta n=\pm 2\mod{5}$. If all the bases in the triplet are assumed to be different from $\{a_{j}\}_{j}$, then it is simple to see that one ends with a similar set of conditions for $n_{X}$, $n_{Y}$ and $n_{Z}$ separating the two sets of triplets as before. Hence, $U^{L}$ maps the elements of $\mathcal{P}_{1}$ in elements of $\mathcal{P}_{1}$ and the elements of $\mathcal{P}_{2}$ in elements of $\mathcal{P}_{2}$.
In the case $M\neq 0,L=0$, we need the following identities,
$\Phi^{\dagger}U\Phi=U\Phi P_{jk}U$ (24)
$\Phi^{\dagger}U^{2}\Phi=U^{3}\Phi U^{2}P^{\prime}_{jk}.$ (25)
Again, we will neglect the contributions of permutations and global phase shifts. From Eq. (20), we have $m_{\overline{X}}=M$ and $n_{\overline{X}}=N$. By exploiting Eqs. (24) and (25), we find that the left members of Eqs. (21) and (22) yield
$U^{n_{\overline{X}}}\Phi^{m_{\overline{X}}}U^{n_{k}}\Phi^{m_{k}}=\delta_{n_{k},0}(1-\delta_{n_{j},0})U^{n_{\overline{X}}}+\delta_{n_{k},\pm 1}$
$(1-\delta_{n_{j},\pm 1})U^{n_{\overline{X}}+n_{k}}\Phi^{m_{k}}+\delta_{n_{k},2}(1-\delta_{n_{j},2})\delta_{m_{\overline{X}},-1}$
$U^{n_{\overline{X}}+n_{k}+1}\Phi^{m_{k}}+\delta_{n_{k},-2}(1-\delta_{n_{j},-2})\delta_{m_{\overline{X}},1}U^{n_{\overline{X}}+n_{k}-1}$
$\Phi^{m_{k}}$ (26)
where $\delta_{a,b}$ is the Kronecker delta, $k\neq j$ and $k,j\in\{Y,Z\}$. By requiring the equality with the right members of Eqs. (21) and (22), we find that the deltas identify the following cases:
  • $n_{\overline{Y}}=n_{\overline{X}}\pm 1\mod{5}\\ n_{\overline{Z}}=n_{\overline{X}}\mp 1\mod{5}$
    implying $\Delta n=n_{Z}-n_{Y}=n_{\overline{Z}}-n_{\overline{Y}}=\pm 2\mod{5}$: the transformations are automorphisms of $\mathcal{P}_{2}$;
  • $n_{\overline{Y}}=n_{\overline{X}}\pm 1\mod{5}\\ n_{\overline{Z}}=n_{\overline{X}}-2\mod{5}$
    implying $\Delta n=n_{\overline{Z}}-n_{\overline{Y}}-1$ and we have that either $\Delta n=1\mod{5}$ or $\Delta n=-2\mod{5}$; in the first case, the transformations are automorphisms of $\mathcal{P}_{1}$, while, in the second one, of $\mathcal{P}_{2}$;
  • $n_{\overline{Y}}=n_{\overline{X}}\pm 1\mod{5}\\ n_{\overline{Z}}=n_{\overline{X}}+2\mod{5}$
    implying $\Delta n=n_{\overline{Z}}-n_{\overline{Y}}+1$ and we have that either $\Delta n=-1\mod{5}$ or $\Delta n=2\mod{5}$; in the first case, the transformations are automorphisms of $\mathcal{P}_{1}$, while, in the second one, of $\mathcal{P}_{2}$;
  • $n_{\overline{Y}}=n_{\overline{X}}-2\mod{5}\\ n_{\overline{Z}}=n_{\overline{X}}+2\mod{5}$
    implying $\Delta n=n_{\overline{Z}}-n_{\overline{Y}}+2=1\mod{5}$: the transformations are automorphisms of $\mathcal{P}_{1}$.
Then, the automorphisms of the whole set of triplets of probability vectors $\mathcal{P}_{1}\cup\mathcal{P}_{2}$ are automorphisms of the disjoint sets $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$. ∎
Note that, while it has been proved [8] that the elements of $\mathcal{S}_{1}$ cannot be mapped into elements of $\mathcal{S}_{2}$ via unitary transformations, Lemma 1 does not imply that there are no unitaries mapping elements of $\mathcal{P}_{1}$ into elements of $\mathcal{P}_{2}$. This is a relevant distinction between the sets $\mathcal{S}_{1}$, $\mathcal{S}_{2}$ and the sets $\mathcal{P}_{1}$, $\mathcal{P}_{2}$. Lemma 1 establish that $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ are distinct equivalence classes, which does not rule out that their elements can be unitarily related. For instance, this is trivially the case if $|\psi_{0}\rangle$ and $|\psi\rangle=V|\psi_{0}\rangle$ are states of MUBs: given two inequivalent triplets of MUBs, one can always pick two states from different MUBs providing the same triplets of probability vectors.
Since the Shannon entropies in the EURs here addressed depend on the Born probabilities only, Lemma 1 is a sufficient condition for the EUR in Eq. (1) (Eq. (2)) to hold for all the triplets in the set $\mathcal{S}_{1}$ ($\mathcal{S}_{2}$). On the other hand, the existence of distinct lower bounds is a necessary, but not sufficient, condition for the inequivalence of $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$, since it implies that there is no unitary transformation that can map the optimal states saturating the EUR for the triplets in $\mathcal{S}_{2}$ into equivalent optimal states for the triplets in $\mathcal{S}_{1}$.